OpenStudy (anonymous):

Unit Circle memorization tips? I have some I see patterns. I can remember COS and Sin (x and Y)...what else y'all got?

OpenStudy (anonymous):

if you remember a one of the sections you can figure out the others fairly easily because they use all the same points

OpenStudy (campbell_st):

think about how it related to the curves sin(a) starts at zero... cos(a) starts at 1. remember the quadrants All, Sine, Tan, Cos

OpenStudy (anonymous):

Also always remember that every angle that is complementary, that is their sum is equal to 90, has the equal recirprocal sine and cosine. For example, 30 + 60 = 90 Sin(30) = Cos(60) = 0.5

OpenStudy (anonymous):

all, students, take, calc

OpenStudy (anonymous):

that's a good one...30+60+90

OpenStudy (anonymous):

Yeah helps alot. Sin(45) = Cos(45) Sin(20) = Cos(70) Cos(20) = Sin(70) Will help you alot. It helped me. And if I have an advice for you, I would say just memorize the most frequent angles, that is 30,45,60. If you know those, you got everything.

OpenStudy (anonymous):

Heard that before.

OpenStudy (anonymous):

the way i remembed the degrees was if its close to a axis then its 30 away. all the others are 15degrees apart

OpenStudy (anonymous):

also, there are three bow ties in the circle

OpenStudy (anonymous):

that's true Ben

OpenStudy (anonymous):

and the middle ones are always +,- sqrt2/2, +,-sqrt2/2 (sign depends on where it is in relation to the x and y axis

OpenStudy (anonymous):

they are all over too minus the 0.90.180.360

OpenStudy (anonymous):

right just dont try to remember more then you need. 0,1, 1/2, 3sqrt2/2, and sqrt2/2 are the only numbers you need to remember. once you have those memorized just put them in the right positions

OpenStudy (anonymous):

OK. cool! good plan.